# The influence of AGN feedback on star formation in red spiral galaxies

### v1

@@ 1 Introduction
The impact of AGNs on galaxy properties is closely tied to structural features such as bars, which are known to drive gas toward galaxy centers and potentially fuel SMBH accretion (e.g., Diaz-Garcia et al., 2020). However, some studies (e.g., Noguchi, 1996; Lesser, 2022) argue that bar instabilities alone are not sufficient to explain quenching in red spirals, reinforcing the role of AGNs as regulators of star formation (Lesser et al. 2025, in preparation). Additionally, LINERs (low-ionization nuclear emission-line regions; Heckman, 1980) are commonly found in the bulges of early-type spiral galaxies, which have a higher fraction of older stars compared to late-type spirals, and can contain low-luminosity AGNs (LLAGN) through shock-ionized gas (e.g., Heckman, 1980; Ferland & Netzer, 1983; Halpern & Steiner, 1983; Terlevich & Melnick, 1985; Ho et al., 2003; Kauffmann et al., 2003).
@@ 2.1 Cluster membership and morphology
Galaxies were selected from a well-defined sample of 115 low-redshift () clusters, 14 of which were observed with the 3.6-meter Canada-France-Hawaii Telescope (CFHT) using the MegaCam detector in the - and -passband (Rude et al., 2020). In addition, - and -band observations of 31 clusters were obtained with the KPNO 0.9-meter telescope using the T2KA CCD detector and the MOSAIC 8K camera (e.g., López-Cruz et al., 2004; Barkhouse et al., 2007, 2009). Supplementary imaging data were also included from the WINGS survey (Fasano et al., 2006; Varela et al., 2009; Valentinuzzi et al., 2011), providing - and -band observations obtained with the 2.5-meter INT and the MPG/ESO 2.2-meter telescopes (Moretti et al., 2014).
Spectroscopic redshifts were obtained from SDSS and NASA/IPAC Extragalactic Database (NED) to confirm cluster membership. Galaxies were included in our sample if their redshifts are within of the cluster velocity dispersion (Yahil & Vidal, 1977) relative to the brightest cluster galaxy (BCG) of the host cluster.
A k-correction was applied to each galaxy based on redshift and color, and was calculated from Chilingarian et al. (2010) and Chilingarian & Zolotukhin (2012). We used for Milky Way extinction corrections (Schlafly & Finkbeiner, 2011), and no correction was made for host galaxy extinction. Only face-on or nearly face-on spirals with obvious spiral arm structures were included to minimize the effect of dust extinction and to preclude the inclusion of edge-on S0 galaxies.
Red spirals were defined as those that are located within of the red-sequence, where is the color dispersion as measured by a Gaussian fit to the color-magnitude distribution (López-Cruz et al., 2004). Blue spirals are those that are found blueward of the red-sequence in the color-magnitude diagram. Full details of galaxy selection are described in Barkhouse et al. (2025, hereafter Paper 1).
@@ 2.3 Chandra and XMM-Newton X-ray data
Publicly available X-ray data from the Chandra X-ray Observatory Source Catalog Release 2 Series (https://doi.org/10.25574/csc2; Evans et al., 2024) and the XMM-Newton X-ray space telescope archive were matched to galaxy positions using a 2′′search radius. When multiple matches were found, the closest match in position was used. Luminosities were calculated from fluxes in the 0.1–10 keV (Chandra) and 0.2–12 keV (XMM-Newton) bands using , where is the luminosity distance, is the observed flux, is the redshift, and (based on Lehmer et al., 2016; Birchall et al., 2022). When a galaxy was matched in both the Chandra and XMM-Newton catalogs, the flux measurement with the smallest uncertainty was retained. Matching our sample to X-ray data yielded the detection of 23 (7.1%) red spirals and 25 (9.2%) blue spirals.
The detection of AGNs using X-ray data is usually based on selecting galaxies that have an X-ray luminosity above a specific threshold. For example, some studies (e.g., Trouille & Barger, 2010; Trouille et al., 2011; Castelló-Mor et al., 2012; Pons & Watson, 2014) used an X-ray luminosity threshold of erg/s. In contrast, Pons (2016) used a low luminosity limit of erg/s based on their analysis of a subset of AGNs from the 12 Micron Galaxy Survey (Brightman & Nandra, 2011).
For our galaxy sample, we found that 44% of blue spiral galaxies have X-ray luminosities erg/s, with 24% having luminosities erg/s. In contrast, 39% of red spirals have luminosities erg/s and 5% have luminosities erg/s. Figure 1 shows the distribution of X-ray luminosity for our sample of blue (left panel) and red (right panel) spiral galaxies.
@@ 3 Results
3.1 Optical classification from emission-line diagnostic diagrams
@@
Optical emission-line diagnostic diagrams are widely used to classify galaxies based on their dominant ionization source. One of the most established method is the use of the BPT diagram, which utilizes the ratios [O III]/H versus [N II]/H to distinguish AGN from star-forming galaxies. This diagnostic diagram is particularly effective because AGNs typically exhibit an elevated [O III]/H ratio due to the presence of high-energy photons (35 eV) needed for [O III] emission, which are more common in AGN continua than in stellar photoionization (Baldwin et al., 1981). Additionally, AGNs often have a stronger [N II]/H ratio compared to normal star-forming galaxies (Veilleux & Osterbrock, 1987).
@@
Figure 2 shows the distribution of galaxies in the standard BPT diagram using the emission-line ratios [O III]/H versus [N II]/H, enabling classification into star-forming, AGN/Seyfert, and LINERs. We note that we use “AGN/Seyfert” in our figures to avoid specific AGN subtype classifications. The black curve and straight line are from Kauffmann et al. (2003), which separates star-forming, AGN/Seyfert, and LINER regions. For the star-forming region, we find that 79.4% of blue galaxies occupy this region, while only 15.5% of red galaxies are found in this area. In the AGN/Seyfert zone, 18.6% of blue galaxies and 64.8% of red spirals are located in this area. Lastly, in the LINER region we find 2% of blue spirals and 19.7% of red galaxies occupy this area. The number and percent of blue and red galaxies that occupy specific regions of the diagram are tabulated in Table 1, along with the results for other emission-line diagnostic diagrams.
@@
To further refine galaxy classification, we employed the Veilleux & Osterbrock diagrams (hereafter VO87 diagrams; see Figures 3 and 4), which depict alternative line ratios: [O III]/H vs. [S II]/H and [O III]/H vs. [O I]/H (Veilleux & Osterbrock, 1987)111[S II] = (6717 + 6731) Å and [O II] = (3726 + 3729) Å. These diagrams are especially useful for separating narrow-line AGNs from H II region-like galaxies and are sensitive to LINERs, which are often associated with weak AGNs or shock excitation (e.g., Heckman, 1980; Terlevich & Melnick, 1985).
@@
For Figure 3, we find 93.4% of blue galaxies are located in the star-forming region, while 3.3% are found each in the AGN/Seyfert and LINER sections. For red spirals, 27.8% are in the star-forming area, while 5.6% are classified as AGN/Seyfert and 66.7% are LINER galaxies. For Figure 4, 80.6% of blue spirals are in the star-forming region, 12.9% are found in the AGN/Seyfert zone, and 6.4% are in the LINER area. For the red spiral population, only 4.8% are star-forming, 11.9% are classified as AGN/Seyfert, and 83.3% are in the LINER region.
@@
We also incorporated newer diagnostics proposed by Mazzolari et al. (2024, 2025) and Backhaus et al. (2025), including [O III] /H vs. [O III] /[O II], and [Ne III]/[O II]. Figures 5 and 6 shows that many red spirals have high [Ne III]/[O II] and [O III]/[O II] ratios, indicative of AGN-like hard ionization. In Figure 5, 0% of blue spirals are found in the “star-forming or AGN/Seyfert” region, while only 2.9% of red spirals are location in this section of the diagram. Conversely, 100% of blue spirals and 97% of red galaxies are in the “AGN/Seyfert-only” section. In Figure 6, 11.1% of blue spirals are in the “star-forming zone,” 11.1% in the “SF or AGN/Seyfert” section, and 77.8% are found in the “AGN/Seyfert-only” region. For red spirals, 3.6% are located in the “star-forming” area, 33.9% in the “SF or AGN/Seyfert” zone, and 66.1% in the “AGN/Seyfert-only” section.
For sources with incomplete or low S/N in classical emission-line ratio diagrams, we adopted WHAN-like extensions; [O III]/[O II] vs. [N II]/H and [O III]/H vs. [N II]/H, shown in Figures 7 and 8 (Cid Fernandes et al., 2010). These plots revealed AGN or LINER-like activity in galaxies misclassified by BPT due to weak lines or dust. In Figure 7, we find that 87.6% of blue spirals occupy the star-forming region of the figure, 8.2% are AGN/Seyfert, and 4.1% are classified as LINERs. For red spirals, 29.3% are in the star-forming area, 32% are in the AGN/Seyfert zone, and 38.7% are located in the LINER region. For Figure 8, 92.6% of blue galaxies are located in the star-forming region, 7.4% are AGN/Seyfert, and 0% are LINER. For red spirals, 24.2% are classified as star-forming, 59.7% are in the AGN/Seyfert region, and 16.1% are LINERs.
To address galaxies with low S/N in H or [N II], we used the WHAN diagram shown in Figure 9, using the equivalent width (EW) of H versus [N II]/H to distinguish between strong/weak AGNs, star-forming, retired, and passive galaxies (see for details; Cid Fernandes et al., 2010, 2011). Several red spirals classified ambiguously in BPT were confirmed as LINERs in WHAN, particularly those with H EW < 3 Å. Specifically for each region, we find for blue spirals that 74.3% are “pure star-forming’,’ 14.9% are “strong AGN,” only 1.4% are “weak AGN,” 9.4% are “retired galaxies,” and 0% are classified as “passive galaxies.” For red spirals, 16% are pure star-forming, 8.5% are strong AGN, 7.5% are weak AGN, 50.9% are retired galaxies, and 17% are passive galaxies.
@@ 3.2 X-ray emission and AGN identification
Among our galaxy sample, 44% of blue spirals and 39% of red spirals exceed erg s-1; 24% of blue and 5% of red spirals have erg s-1. This suggests AGNs are not exclusive to quiescent systems, supporting the dual-mode feedback model (e.g., Cresci & Maiolino, 2018; Venturi et al., 2023). Only 35% of red and 36% of blue spirals had measurable optical emission lines (see Figures 2 to 9), and most were LINERs or LLAGNs, with weak ionization features (Kauffmann et al., 2003; Silverman et al., 2008). Tables 4 and 5 give the classification of galaxies based on X-ray luminosity and emission-line ratios.
We also calculated the X-ray mass-to-light ratio using extinction- and k-corrected -band absolute magnitudes as a proxy for galaxy mass (Mahajan et al., 2018). We find that blue spirals have an average and red spirals have .

### v2

@@ 1 Introduction
The impact of AGNs on galaxy properties is closely tied to structural features such as bars, which are known to drive gas toward galaxy centers and potentially fuel SMBH accretion (e.g., Diaz-Garcia et al., 2020). However, some studies (e.g., Noguchi, 1996; Lesser, 2022) argue that bar instabilities alone are not sufficient to explain quenching in red spirals, reinforcing the role of AGNs as regulators of star formation (Lesser et al. 2026, in preparation). Additionally, LINERs (low-ionization nuclear emission-line regions; Heckman, 1980) are commonly found in the bulges of early-type spiral galaxies, which have a higher fraction of older stars compared to late-type spirals, and can contain low-luminosity AGNs (LLAGN) through shock-ionized gas (e.g., Heckman, 1980; Ferland & Netzer, 1983; Halpern & Steiner, 1983; Terlevich & Melnick, 1985; Ho et al., 2003; Kauffmann et al., 2003).
@@ 2.1 Cluster membership and morphology
Galaxies were selected from a well-defined sample of 115 low-redshift () clusters, 14 of which were observed with the 3.6-meter Canada-France-Hawaii Telescope (CFHT) using the MegaCam detector in the - and -passband, with data reduction, object detection, and galaxy photometry fully described in Rude et al. (2020). In addition, - and -band observations of 31 clusters were obtained with the KPNO 0.9-meter telescope using the T2KA CCD detector and the MOSAIC 8K camera (see López-Cruz et al., 2004; Barkhouse et al., 2007, 2009, for details regarding object detection and photometry). Supplementary archival imaging data and catalogs were included from the WINGS survey (Fasano et al., 2006; Varela et al., 2009; Valentinuzzi et al., 2011), providing - and -band observations of 70 clusters obtained with the 2.5-meter INT and the MPG/ESO 2.2-meter telescopes (see Moretti et al., 2014, for details regarding catalog and image access).
All galaxy cluster catalogs from the 115-cluster sample have been analyzed in a homogeneous fashion using the same procedure to measure red-sequence slopes, y-intercepts, and dispersions from cluster color-magnitude diagrams (see López-Cruz et al., 2004, for details regarding the fitting method). The faint-end magnitude cut-off for each cluster is based on the limiting magnitude in the //-band, where the mean color uncertainty in the cluster red-sequence exceeds the dispersion of the red-sequence. The red-sequence dispersion () is measured by a Gaussian fit to the color-magnitude distribution (López-Cruz et al., 2004; Barkhouse et al., 2007). All galaxies fainter than the limiting magnitude of the host cluster are excluded from further analysis.
To minimize selection effects, red spirals were defined as those that are located within of the red-sequence. Blue spirals are those that are found blueward of the red-sequence in the color-magnitude diagram. Spectroscopic redshifts were obtained from SDSS and NASA/IPAC Extragalactic Database (NED) to confirm cluster membership and minimize inclusion of interlopers. Galaxies were included in our sample if their redshifts are within of the cluster velocity dispersion (Yahil & Vidal, 1977) relative to the brightest cluster galaxy (BCG) of the host cluster.
A k-correction was applied to each galaxy based on redshift and color, and was calculated from Chilingarian et al. (2010) and Chilingarian & Zolotukhin (2012). We used for Milky Way extinction corrections (Schlafly & Finkbeiner, 2011), and no correction was made for host galaxy extinction. Only face-on or nearly face-on spirals with obvious spiral arm structures were included to minimize the effect of dust extinction and to preclude the inclusion of edge-on S0 galaxies. Spiral galaxies were selected via visual verification to avoid the inclusion of non-spiral systems.
Our final sample of face-on cluster spiral galaxies consists of 273 blue spirals and 324 red spirals. Based on SDSS DR 17 photometry, extinction- and k-corrected absolute magnitudes of blue spirals range from , with an average of , while red spirals range from , with a mean of . The average redshift of our sample of blue spirals is , while the average for red spiral galaxies is .
Using the relationship between and galaxy mass given by Mahajan et al. (2018, see their Figure 2), the mass of blue spirals range from to . The mass range of red spirals extends from to .
@@ 2.3 Chandra and XMM-Newton X-ray data
Publicly available X-ray data from the Chandra X-ray Observatory Source Catalog Release 2 Series (https://doi.org/10.25574/csc2; Evans et al., 2024) and the XMM-Newton X-ray space telescope archive were matched to galaxy positions using a 2′′search radius. When multiple matches were found, the closest match in position was used. Luminosities were calculated from fluxes in the 0.1–10 keV (Chandra) and 0.2–12 keV (XMM-Newton) bands using , where is the luminosity distance, is the observed flux, is the redshift, and (based on Lehmer et al., 2016; Birchall et al., 2022). When a galaxy was matched in both Chandra and XMM-Newton catalogs, the flux measurement with the smallest uncertainty was retained. This occurred for 12 out of 45 galaxies, with a difference in flux exceeding for two of the 12 galaxies ( and , see Appendix A, Figure 1).
Matching our sample to X-ray data yielded a detection for 23 (7.1%) red spirals and 22 (8%) blue spirals. Non-detections of the remaining spirals (552 out of 597, 92.5%) was due to both a lack of coverage by Chandra and XMM-Newton telescopes (239 spirals out of 597, 40%), and coverage but no detection due to sensitivity limits (313 out of 597, 52.4%) . For red spirals, X-ray luminosities range from erg s-1 to erg s-1, with a mean of erg s-1 and a dispersion (rms) of erg s-1. For blue spiral galaxies, X-ray luminosities range from erg s-1 to erg s-1, with an average of erg s-1 and a dispersion (rms) of erg s-1.
The detection of AGNs using X-ray data is usually based on selecting galaxies that have an X-ray luminosity above a specific threshold. For example, some studies (e.g., Trouille & Barger, 2010; Trouille et al., 2011; Castelló-Mor et al., 2012; Pons & Watson, 2014) used an X-ray luminosity threshold of erg s-1. In contrast, Pons (2016) used a low luminosity limit of erg s-1 based on their analysis of a subset of AGNs from the 12 Micron Galaxy Survey (Brightman & Nandra, 2011).
For our X-ray-detected galaxy sample, we found that 9 out of 22 blue spirals (40.9%) have X-ray luminosities erg s-1, with 22.7% (5 out of 22) having luminosities erg s-1. For the red spirals, 39.1% (9 out of 23) have luminosities erg s-1 and 4.3% (1 out of 23) have luminosities erg s-1. Figure 1 shows the distribution of X-ray luminosity for our sample of blue and red spiral galaxies.
@@ 3 Results
3.1 [O III]/H vs [N II]/H
@@
Optical emission-line diagnostic diagrams are widely used to classify galaxies based on their dominant ionization source. One of the most established method is the use of the BPT diagram, which utilizes the ratios [O III]/H versus [N II]/H to distinguish AGN from star-forming galaxies. This diagnostic diagram is particularly effective because AGNs typically exhibit an elevated [O III]/H ratio due to the presence of high-energy photons (35 eV) needed for [O III] emission, which are more common in AGN continua than in stellar photoionization (Baldwin et al., 1981).
Additionally, AGNs often have a stronger [N II]/H ratio compared to normal star-forming galaxies (Veilleux & Osterbrock, 1987). The physical origin of this enhancement lies in the hardness of the AGN radiation field. While young O- and B-type stars in star-forming galaxies produce a relatively soft ionizing spectrum that is rapidly absorbed, AGNs generate a hard power-law continuum extending into the extreme-UV and X-ray regime (Baldwin et al., 1981; Ho, 2008). These high-energy photons penetrate deep into the surrounding interstellar medium, creating extended partially ionized zones where neutral hydrogen and N+ ions coexist (Ferland & Netzer, 1983; Kewley et al., 2019). In such regions, energetic electrons efficiently collisionally excite nitrogen ions, significantly boosting the [N II] emission relative to the recombination-dominated H line. This effect can be further enhanced by shocks driven by AGN outflows or jets, which heat and compress the gas and increase the efficiency of low-ionization line emission (Heckman, 1980; Rich et al., 2011; Farage et al., 2010). In contrast, H II regions dominated by stellar photoionization have much thinner partially ionized zones and lack strong shock excitation, resulting in the characteristically lower [N II]/H ratios observed in star-forming galaxies.
@@
Figure 2 shows the distribution of galaxies in the standard BPT diagram using the emission-line ratios [O III]/H versus [N II]/H, enabling classification into star-forming, AGN/Seyfert, and LINERs. We note that we use “AGN/Seyfert” in our figures to avoid specific AGN subtype classifications. The black curve and straight line are from Kauffmann et al. (2003), which separates star-forming, AGN/Seyfert, and LINER regions. For the star-forming region, we find that 79.4% of blue galaxies (81 out of 102) occupy this region, while only 15.5% of red galaxies (11 out of 71) are found in this area. In the AGN/Seyfert zone, 18.6% (19) of blue galaxies and 64.8% (46) of red spirals are located in this area. Lastly, in the LINER region we find 2% (2) of blue spirals and 19.7% (14) of red galaxies occupy this area. The number and percent of blue and red galaxies that occupy specific regions of the diagram are tabulated in Table 1, along with the results for other emission-line diagnostic diagrams.
3.2 [O III]/H vs. [S II]/H
@@
To further refine galaxy classification, we employed the Veilleux & Osterbrock diagrams (hereafter VO87 diagrams; see, for example, Figure 3), which depict alternative line ratios: [O III]/H vs. [S II]/H (Veilleux & Osterbrock, 1987)111[S II] = (6717 + 6731) Å and [O II] = (3726 + 3729) Å. This diagram is especially useful for separating narrow-line AGNs from H II region-like galaxies and are sensitive to LINERs, which are often associated with weak AGNs or shock excitation (e.g., Heckman, 1980; Terlevich & Melnick, 1985).
@@
For Figure 3, we find 93.4% (85 out of 91) of blue galaxies are located in the star-forming region, while 3.3% (3) are found each in the AGN/Seyfert and LINER sections. For red spirals, 27.8% (15 out of 54) are in the star-forming area, while 5.6% (3) are classified as AGN/Seyfert and 66.7% (36) are LINER galaxies.
3.3 [O III]/H vs. [O I]/H
In Figure 4, we use the VO97 diagram of [O III]/H vs. [O I]/H to help separate AGN/Seyfert galaxies from star-forming and LINER-like systems. For Figure 4, we find that 25 out of 31 blue spirals (80.6%) are located in the star-forming region, 12.9% (4) are found in the AGN/Seyfert zone, and 6.4% (2) are in the LINER area. For the red spiral population, only 4.8% (2 out of 42) are star-forming, 11.9% (5) are classified as AGN/Seyfert, and 83.3% (35) are in the LINER region.
@@
3.4 [O III] /H vs. [O III] /[O II]
We incorporated newer diagnostics proposed by Mazzolari et al. (2024, 2025) and Backhaus et al. (2025), including [O III] /H vs. [O III] /[O II]. Figure 5 shows that many red spirals have high [Ne III]/[O II] and [O III]/[O II] ratios, indicative of AGN-like hard ionization. In Figure 5, 0% of blue spirals (0 out of 10) are found in the “star-forming or AGN/Seyfert” region, while only 2.9% (2 out of 68) of red spirals are location in this section of the diagram. Conversely, 100% (10) of blue spirals and 97% (66) of red galaxies are in the “AGN/Seyfert-only” section.
3.5 [O III] /H vs. [Ne III]/[O II]
An additional diagnostic plot proposed by Mazzolari et al. (2024, 2025) and Backhaus et al. (2025) is the use of [O III] /H vs. [Ne III]/[O II] to help discriminate between star-forming galaxies and AGN/Seyfert systems. Our data are plotted in Figure 6, where 11.1% of blue spirals (1 out of 9) are in the “star-forming” zone, 11.1% (1) in the “SF or AGN/Seyfert” section, and 77.8% (7) are found in the “AGN/Seyfert-only” region. For red spirals, 3.6% (2 out of 56 galaxies) are located in the “star-forming” area, 30.3% (17) in the “SF or AGN/Seyfert” zone, and 66.1% (37) in the “AGN/Seyfert-only” section.
3.6 [O III]/[O II] vs. [N II]/H
For sources with incomplete or low S/N in classical emission-line ratio diagrams, we adopted the WHAN-like222The WHAN diagram, introduced by Cid Fernandes et al. (2010, 2011), combines the H equivalent width with the [N II]/H ratio to classify galaxies and to include weak-line systems that are excluded from classical BPT diagrams due to low H or [O III] signal-to-noise. extension [O III]/[O II] vs. [N II]/H as shown in Figure 7 (Cid Fernandes et al., 2010). This plot reveals AGN or LINER-like activity in galaxies misclassified by BPT due to weak lines or dust. In Figure 7, we find that 85 of 97 blue spirals (87.6%) occupy the star-forming region of the figure, 8.2% (8) are AGN/Seyfert, and 4.1% (4) are classified as LINERs. For red spirals, 29.3% (22 of 75 galaxies) are in the star-forming area, 32% (24) are in the AGN/Seyfert zone, and 38.7% (29) are located in the LINER region.
3.7 [O III]/H vs. [N II]/H
An additional WHAN-like diagnostic plot uses [O III]/H vs. [N II]/H to help separate star-forming, AGN/Seyfert, and LINER systems (see Figure 8). Of the 95 blue galaxies plotted in Figure 8, 92.6% (88 out of 95) are located in the star-forming region, 7.4% (7) are AGN/Seyfert, and 0% are LINER. For 62 red spirals, 24.2% (15) are classified as star-forming, 59.7% (37) are in the AGN/Seyfert region, and 16.1% (10) are LINERs.
3.8 Equivalent width of H versus [N II]/H
To address galaxies with low S/N in H or [N II], we used the WHAN diagram shown in Figure 9, using the equivalent width (EW) of H versus [N II]/H, to distinguish between strong/weak AGNs, star-forming, retired, and passive galaxies (see for details; Cid Fernandes et al., 2010, 2011). Several red spirals classified ambiguously in BPT were confirmed as LINERs in WHAN, particularly those with H EW < 3 Å. Of the 155 blue spiral galaxies, 74.8% (116) are “pure star-forming,” 14.2% (22) are “strong AGN,” only 1.3% (2) are “weak AGN,” 9.0% (14) are “retired galaxies,” and 0.6% (1) are classified as “passive galaxies.” Of the 114 red spiral galaxies, 15.8% (18) are pure star-forming, 7.9% (9) are strong AGN, 7.0% (8) are weak AGN, 41.2% (47) are retired galaxies, and 28.1% (32) are passive galaxies.
@@ 3.9 X-ray emission and AGN identification
Among our X-ray-detected galaxy sample, 40.9% of blue spirals (9 out of 22) and 39.1% of red spirals (9 out of 23) exceed erg s-1; 22.7% of blue (5 out of 22) and 4.3% of red spirals (1 out of 23) have erg s-1. This suggests AGNs are not exclusive to quiescent systems, supporting the dual-mode feedback model (e.g., Cresci & Maiolino, 2018; Venturi et al., 2023). In this framework, AGNs operate in two distinct modes based on the energy transfer mechanism. The higher-luminosity AGNs in our blue spiral sample correspond to the “radiative mode” (or quasar mode), where energy is released via radiation or winds from the accretion disk, ionizing and heating the surrounding gas (Morganti, 2017). This process regulates star formation by preventing gas from cooling or by expelling it through quasar-driven winds (Harrison et al., 2014; Rubinur, 2024). Conversely, the X-ray activity in our red spirals likely represents the “kinetic mode” (or radio mode), dominant in lower-power AGNs where mechanical energy is injected into the medium via jets and shocks (Morganti, 2017). The presence of significant X-ray luminosities across both populations indicates that AGN feedback is a multi-stage process, transitioning from radiative-dominated quenching in gas-rich blue systems to a kinetic-dominated mechanism in quiescent red systems.
Only 35% (113 out of 324) red spirals and 36% (98 out of 273) blue spirals had measurable optical emission lines (see Figures 2 to 9), and most were LINERs or LLAGNs, with weak ionization features (Kauffmann et al., 2003; Silverman et al., 2008). The discrepancy between our X-ray detection rates and optical line measurements suggests that a significant fraction of AGN activity in these spirals is obscured or diluted by host galaxy light, particularly in massive, star-forming systems. Tables 4 and 5 gives the classification of galaxies based on X-ray luminosity and emission-line ratios.